OpenStudy (anonymous):

Solve 3(x – 5) + 5(4 – x) = 8. Select one: a. x = -3/2 b. x = -13/2 c. x = 13/8 d. x = 3/8

OpenStudy (anonymous):

I'd say a. but i've got 3/2 and not-3/2.

OpenStudy (yanasidlinskiy):

Distributive Property: 3x-15+20-5x = 8 Combine like terms on both sides -2x+5 = 8 Subtract the number that is added to the x. Always do the opposite. -5 -5 -2x = 3 Divide by -2 x = -0.6

OpenStudy (anonymous):

Or yeah, that's right :-) a.

OpenStudy (yanasidlinskiy):

Which is the closest to a. Like @Sulle26 said.

OpenStudy (anonymous):

oh alright thanks guys.(:

OpenStudy (yanasidlinskiy):

No problem!:)

OpenStudy (yanasidlinskiy):

Glad to help!:)

OpenStudy (anonymous):

no problemo :)